Given an activity's optimistic, most likely, and pessimistic time estimates of 5, 7, and 15 days respectively, compute the PERT expected activity time (t) and variance of activity completion time (Var) for this activity.

1. t = 8, and Var = 2.78

2. t = 7, and Var = 1.78

3. t = 14, and Var = 2.78

4. t = 8, and Var = 1.78
Answer #1

+ [to + 4 dent to 5+4 (1) + 15) 11 [5 +28+150 [48] = 18/6=2 variance face tp-to]?? Lucis-s)] 100 V 2-78 36 Carred aption it =